OpenStudy (anonymous):

What is the Value of X/Y when X = 9/4 and Y =3/5? A.15/4 B.4/3 C.27/20 D.- 15/4

OpenStudy (anonymous):

c.

OpenStudy (anonymous):

@Clalgee

OpenStudy (anonymous):

@MathDummy-123

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@iPwnBunnies

zepdrix (zepdrix):

\[\Large\rm \color{orangered}{X=\frac{9}{4}},\qquad\qquad \color{royalblue}{Y=\frac{3}{5}}\] \[\Large\rm \frac{\color{orangered}{X}}{\color{royalblue}{Y}}\quad=\quad \frac{\color{orangered}{9/4}}{\color{royalblue}{3/5}}\]Remember what to do when you `divide` by a fraction?

zepdrix (zepdrix):

You flip the bottom fraction, and rewrite the operation as multiplication. Sound familiar?

zepdrix (zepdrix):

\[\Large\rm \frac{9}{4}\cdot\frac{5}{3}\]Then just multiply across, yes? Understand?

OpenStudy (anonymous):

OK I'm correct C. Its 9x3 4x5

zepdrix (zepdrix):

No no no no no no no. I see this mistake so often.

zepdrix (zepdrix):

You only cross multiply if there is an `equals sign`. Normal multiplication is just top multiply top, bottom multiply bottom.

zepdrix (zepdrix):

\[\Large\rm \frac{9}{4}=\frac{5}{3}\qquad\to\qquad 9\cdot3=4\cdot5\]\[\Large\rm \frac{9}{4}\cdot\frac{5}{3}=\frac{9\cdot5}{4\cdot3}\]

OpenStudy (anonymous):

Ok still dont get the answer

zepdrix (zepdrix):

Because you have to simplify.

OpenStudy (jdoe0001):

\(\bf \cfrac{x}{y}\qquad x=\frac{9}{4}\qquad y=\frac{3}{5} \\ \quad \\ \cfrac{x}{y}\implies \cfrac{\frac{9}{4}}{\frac{3}{{\color{brown}{ 5}}}}\implies \cfrac{9}{4}\cdot \cfrac{{\color{brown}{ 5}}}{3}\implies \cfrac{\cancel{ 9 }\cdot 5}{4\cdot \cancel{ 3 }}\implies ?\)